# Attention as an Adaptive Filter

### v1

@@ 3.8.1 Adaptive Filter Attention
We now present an implementation of the above algorithm, for the case of isotropic noise and decay.
@@
We switch from the rational weight $\tfrac{1}{1+r_{ij}^{2}/\nu}$ to the equivalent exponential–log form $\exp\!\big{(}-\tau_{s}\log(1+r_{ij}^{2}/\nu)\big{)}$, which both improves numerical stability and introduces a temperature parameter $\tau_{s}$ that controls the sharpness of the weighting; this reduces exactly to the rational form when $\tau_{s}=1$. We drop $\lambda_{C}^{2}$ since we can absorb it into $\sigma^{2}$, and drop the Mahalanobis scale $\nu$, since we can absorb it into $\sigma,\eta,\tau_{s},\mathbf{W}_{Q},\mathbf{W}_{K}$.
@@
Another option to enforce real inputs is to impose a penalty on $\|\operatorname{Im}(\mathbf{Z_{p}})\|$ or on $\|\mathbf{W_{v}}\mathbf{W_{p}}-\mathbf{I}^{d_{v}\times d_{v}}\|$. For simplicity, we do not impose penalties and allow these matrices to freely vary.
@@
(Define $t_{i}:=\Delta t\cdot i,\quad t_{j}:=\Delta t\cdot j\$.) Complex-valued Projections:
@@
$\|\mathbf{R}_{\mathbf{qk}}\|^{2}=\mathbf{1}_{m}(\mathbf{1}_{d}^{\top}\|\mathbf{Z_{q}}\|^{2})+\mathbf{E_{qk}}^{2}\odot(\mathbf{1}_{d}^{\top}\|\mathbf{Z_{k}}\|^{2})^{\top}\mathbf{1}_{m}^{\top}-2\mathbf{E_{qk}}\odot\mathrm{Re}(\mathbf{\tilde{Z}}_{\mathbf{q}}^{\dagger}\mathbf{\tilde{Z}}_{\mathbf{k}})$
Attention matrix:
$\mathbf{L}=-\log\left(\mathbf{V}^{C}+\|\mathbf{R}_{\mathbf{qk}}\|^{2}+\epsilon\right)\odot\mathbf{M}_{\text{causal}},$
@@
Output with value dynamics: $\mathbf{\bar{Z}_{v}}=\bm{\tilde{\Phi}}_{\mathbf{v}}^{+}\odot(\mathbf{\tilde{Z}}_{\mathbf{v}}\mathbf{\hat{A}}^{\dagger})$
@@
Prediction: $\mathbf{Z}_{\mathbf{v}}^{\text{pred}}=(e^{\alpha\Delta t}e^{\bm{\lambda}_{\Omega}\Delta t})\odot\mathbf{\bar{Z}_{v}}$
Output Projection: $\big{(}\operatorname{Re}(\mathbf{Z_{p}}),\operatorname{Im}(\mathbf{Z_{p}})\big{)}\leftarrow\mathcal{L}_{P}\big{(}\operatorname{Re}(\mathbf{Z}_{\mathbf{v}}^{\text{pred}}),\operatorname{Im}(\mathbf{Z}_{\mathbf{v}}^{\text{pred}})\big{)}$
@@
Algorithm 1 Adaptive Filter Attention (Single Head)
@@
3.8.2 Unrolled Adaptive Filter Attention
Complex-valued Projections:
State transition matrix:
for $i\leftarrow 1$ to $m$ do
 $p_{i}^{-(0)}=1/(\Phi_{\alpha}^{2}/p_{i-1}^{+}+\sigma^{2})$
    // Propagated precision
 for $k\leftarrow 0$ to K1K-1
   // Inner iteration within time step
   do
    Recompute queries and keys:
    Compute innovations: $\mathbf{y}_{\mathbf{qk}i}^{(k)}=\mathbf{z}_{\mathbf{q}i}-\mathbf{\bar{z}}_{\mathbf{k}i}^{-(k)},\quad\mathbf{y}_{\mathbf{v}i}^{(k)}=\mathbf{z}_{\mathbf{v}i}-\mathbf{\bar{z}}_{\mathbf{v}i}^{-(k)}$
    Posterior mean update:
 Propagate mean and precision to next time step:
Return: $\operatorname{Re}(\mathbf{z_{p}})$
Algorithm 2 Unrolled Adaptive Filter Attention (Inference)
@@ 3.9 Basic Experiments
While standard attention is also able to filter this noisy trajectory, it is not able to do so effectively in a single layer, and the resulting attention matrices lack the consistent and interpretable attention matrices of AFA. We compare the attention matrices from a trained model comprised of two layers of standard attention to a trained model comprised of a single layer of AFA in Fig. 3. The AFA attention matrix displays a clear periodic structure, corresponding to pair-wise precisions propagated through the underlying oscillatory dynamics.
@@ 4 Radial-Tangential Model
We were able to significantly simplify our algorithm by assuming that the dynamics decouple into separate magnitude and angular components, and that the process and measurement noise are isotropic. Here, we will slightly relax these assumptions by assuming that the process and measurement noise also decompose into separate radial and tangential components, and derive a modification of the above algorithm.
@@ 4.1 Itô SDE with Radial–Tangential Dynamics and Noise
$d\mathbf{x}(t)=\mathbf{A}\mathbf{x}(t)dt+\mathbf{B}(t)d\mathbf{w}(t),\quad\mathbf{z}(t_{k})=\mathbf{C}\mathbf{x}(t_{k})+\mathbf{v}(t_{k}),$
Let $\mathbf{A}(t)=\mathbf{S}\bm{\Lambda}\mathbf{S}^{-1}$, with $\bm{\Lambda}=\alpha\mathbf{I}+\bm{\Lambda}_{\Omega}$, where $\bm{\Lambda}_{\Omega}\in i\mathbb{R}^{d\times d}$ is diagonal, and $\mathbf{B}(t)=\mathbf{S}\bm{\Lambda}_{B}(t)\mathbf{S}^{\dagger}$, $\mathbf{C}(t)=\lambda_{C}\mathbf{I}$. We can also write the eigenvalues as $\bm{\lambda}=\alpha\mathbf{1}+\bm{\omega}$, where $\alpha\in\mathbb{R}$, $\bm{\omega}=\text{diag}(\bm{\Lambda}_{\Omega})\in i\mathbb{R}^{d}$.
@@
$m_{\mathbf{s}}(t)=\|\mathbf{x}_{\mathbf{s}}(t)\|_{2}=\sqrt{\mathbf{x}_{\mathbf{s}}^{\dagger}(t)\mathbf{x}_{\mathbf{s}}(t)}\in\mathbb{R},\quad\mathbf{u}_{\mathbf{s}}(t)=\mathbf{x}_{\mathbf{s}}(t)/\|\mathbf{x}_{\mathbf{s}}(t)\|_{2}\in\mathbb{S}^{d-1}.$
$m_{\mathbf{zs}}(t)=\|\mathbf{z}_{\mathbf{s}}(t)\|_{2}=\sqrt{\mathbf{z}_{\mathbf{s}}^{\dagger}(t)\mathbf{z}_{\mathbf{s}}(t)}\in\mathbb{R},\quad\mathbf{u}_{\mathbf{zs}}(t)=\mathbf{z}_{\mathbf{s}}(t)/\|\mathbf{z}_{\mathbf{s}}(t)\|_{2}\in\mathbb{S}^{d-1}.$
@@
$\exp_{\mathbf{u}}:T_{\mathbf{u}}\mathbb{S}^{d-1}\to\mathbb{S}^{d-1},\quad\exp_{\mathbf{u}}(\mathbf{v})=\cos\big(\|\mathbf{v}\|\big{missing})\,\mathbf{u}+\sin\big(\|\mathbf{v}\|\big{missing})\,\frac{\mathbf{v}}{\|\mathbf{v}\|},$
@@ 4.2 Propagation of Uncertainty through the RT-SDE
$\mathbf{x}(t_{i})=e^{\mathbf{A}(t_{i}-t_{j})}\mathbf{x}(t_{j})+\int_{t_{j}}^{t_{i}}e^{\mathbf{A}(t_{i}-s)}\mathbf{B}(s)d\mathbf{w}(s)$
@@
$\mathbf{x}(t_{i})=e^{-\mathbf{A}\Delta t_{ij}}\left(\mathbf{x}(t_{j})+\int_{0}^{-\Delta t_{ij}}e^{-\mathbf{A}\tau}\mathbf{B}(t_{j}+\tau)d\mathbf{w}(\tau)\right).$
@@
$\bm{\Lambda}_{\mathbf{V}}^{C}(t_{i},t_{j})=\lambda_{C}^{2}\bm{\Lambda}_{\mathbf{V}F}(t_{i},t_{j})+e^{-\bm{\Lambda}\Delta t_{ij}}\bm{\Lambda}_{R}(t_{j})e^{-\bm{\Lambda}^{\dagger}\Delta t_{ij}}$
$\lambda_{C}^{2}\bm{\Lambda}_{\mathbf{V}F}(t_{i},t_{j})+e^{-\bm{\Lambda}\Delta t_{ij}}\bm{\Lambda}_{R}(t_{j})e^{-\bm{\Lambda}^{\dagger}\Delta t_{ij}}$
$\lambda_{C}^{2}\bm{\Lambda}_{\mathbf{V}F}(t_{i},t_{j})+e^{-2\alpha\Delta t_{ij}}\bm{\Lambda}_{R}(t_{i})$
@@
$\bm{\Lambda}_{\mathbf{V}}^{C}(t_{i},t_{j})$ $=\Big{(}\lambda_{C}^{2}\kappa_{F}(|\Delta t_{ij}|)\sigma_{t}^{2}+e^{-2\alpha\Delta t_{ij}}\eta_{t}^{2}\Big{)}\mathbf{I}$
$\quad+\Big{(}\lambda_{C}^{2}\kappa_{F}(|\Delta t_{ij}|)(\sigma_{r}^{2}-\sigma_{t}^{2})+e^{-2\alpha\Delta t_{ij}}(\eta_{r}^{2}-\eta_{t}^{2})\Big{)}\mathbf{u_{s}}(t_{i})\mathbf{u_{s}}(t_{i})^{\dagger}$
@@
$\sigma_{Vr}^{2}(|\Delta t_{ij}|)$ $=\lambda_{C}^{2}\,\kappa_{F}(|\Delta t_{ij}|)\,\sigma_{r}^{2}+e^{-2\alpha\Delta t_{ij}}\,\eta_{r}^{2},$
$\sigma_{Vt}^{2}(|\Delta t_{ij}|)$ $=\lambda_{C}^{2}\,\kappa_{F}(|\Delta t_{ij}|)\,\sigma_{t}^{2}+e^{-2\alpha\Delta t_{ij}}\,\eta_{t}^{2}.$
@@
$\bm{\Lambda}_{\mathbf{V}}^{C}(t_{i},t_{j})=\Big{(}\lambda_{C}^{2}\kappa_{F}(|\Delta t_{ij}|)\sigma^{2}+e^{-2\alpha\Delta t_{ij}}\eta^{2}\Big{)}\mathbf{I}$
@@ 4.3 Radial–Tangential Kalman IRLS
$m_{\mathbf{zs}i}^{(k)}=\|\mathbf{z}_{\mathbf{s}i}^{(k)}\|_{2},\quad\mathbf{u}_{\mathbf{zs}i}^{(k)}=\mathbf{z}_{\mathbf{s}i}^{(k)}/\|\mathbf{z}_{\mathbf{s}i}^{(k)}\|_{2}.$
@@
$\begin{gathered}\bm{\Lambda}_{\mathbf{V}}^{C}(t_{i},t_{j})=\sigma_{Vt}^{2}(|\Delta t_{ij}|)\big{(}\mathbf{I}-\mathbf{u_{s}}(t_{i})\mathbf{u_{s}}(t_{i})^{\dagger}\big{)}+\sigma_{Vr}^{2}(|\Delta t_{ij}|)\mathbf{u_{s}}(t_{i})\mathbf{u_{s}}(t_{i})^{\dagger},\\[5.69054pt] \sigma_{Vr}^{2}(|\Delta t_{ij}|)=\lambda_{C}^{2}\kappa_{F}(|\Delta t_{ij}|)\sigma_{r}^{2}+e^{-2\alpha\Delta t_{ij}}\eta_{r}^{2},\\[2.84526pt] \sigma_{Vt}^{2}(|\Delta t_{ij}|)=\lambda_{C}^{2}\kappa_{F}(|\Delta t_{ij}|)\sigma_{t}^{2}+e^{-2\alpha\Delta t_{ij}}\eta_{t}^{2}.\end{gathered}$
@@
$\mathbf{P}_{T\mathbf{zs}i}^{(k)}\Big{(}\mathbf{z}_{\mathbf{s}i}^{(k)}/\lambda_{C}\Big{)}:=\mathbf{I}-\frac{\mathbf{z}_{\mathbf{s}i}^{(k)}\Big{(}\mathbf{z}_{\mathbf{s}i}^{(k)}\Big{)}^{\dagger}/\lambda_{C}^{2}}{\Big{(}\mathbf{z}_{\mathbf{s}i}^{(k)}\Big{)}^{\dagger}\Big{(}\mathbf{z}_{\mathbf{s}i}^{(k)}\Big{)}/\lambda_{C}^{2}}=\mathbf{I}-\mathbf{u}_{\mathbf{zs}i}^{(k)}(\mathbf{u}_{\mathbf{zs}i}^{(k)})^{\dagger}.$
@@
$\theta_{i}^{(k)}:=\big{\|}\mathbf{\bar{v}}_{\mathbf{zs}i}^{(k)}\big{\|},\qquad\mathbf{a}_{i}^{(k)}:=\frac{\mathbf{\bar{v}}_{\mathbf{zs}i}^{(k)}}{\|\mathbf{\bar{v}}_{\mathbf{zs}i}^{(k)}\|},$
@@
$\mathbf{y}_{\mathbf{zs}i}^{(k+1)}=\mathbf{u}_{\mathbf{zs}i}^{(k)}+\mathbf{\bar{v}}_{\mathbf{zs}i}^{(k)}/m_{\mathbf{zs}i}^{(k)},\qquad\mathbf{u}_{\mathbf{zs}i}^{(k+1)}=\frac{\mathbf{y}_{\mathbf{zs}i}^{(k+1)}}{\|\mathbf{y}_{\mathbf{zs}i}^{(k+1)}\|},$
@@
$\mathbf{y}_{\mathbf{zs}i}^{(k+1)}=\mathbf{z}_{\mathbf{s}i}^{(k)}+\mathbf{\bar{v}}_{\mathbf{zs}i}^{(k)},\qquad\mathbf{u}_{\mathbf{zs}i}^{(k+1)}=\frac{\mathbf{y}_{\mathbf{zs}i}^{(k+1)}}{\|\mathbf{y}_{\mathbf{zs}i}^{(k+1)}\|},$

### v2

@@ 4.1 Algorithm
We now present an implementation of Adaptive Filter Attention, for the case of isotropic noise and decay.
@@
We switch from the rational weight $\tfrac{1}{1+r_{ij}^{2}/\nu}$ to the equivalent exponential–log form $\exp\!\big(-\tau_{s}\log(1+r_{ij}^{2}/\nu)\big)$, where $\tau_{s}$ is the usual temperature parameter that controls the sharpness of the weighting, with the aim of improving numerical stability; this reduces exactly to the rational form when $\tau_{s}=1$. We drop $\big|\lambda_{C}\big|^{2}$ since we can absorb it into $\sigma^{2}$, and drop the Mahalanobis scale $\nu$, since we can absorb it into $\sigma,\eta,\tau_{s},\mathbf{W}_{Q},\mathbf{W}_{K}$.
@@
Another option to enforce real inputs is to impose a penalty on $\big|\operatorname{Im}(\mathbf{Z_{p}})\big|$ or on $\big|\mathbf{W_{v}}\mathbf{W_{p}}-\mathbf{I}^{d_{v}\times d_{v}}\big|$. For simplicity, we do not impose penalties and allow these matrices to freely vary.
@@
(Define $t_{i}:=\Delta t\cdot i,\quad t_{j}:=\Delta t\cdot j\$.)
1exComplex-valued Projections:
@@
$\big|\mathbf{R}_{\mathbf{qk}}\big|^{2}=\mathbf{1}_{m}(\mathbf{1}_{d}^{\top}\big|\mathbf{Z_{q}}\big|^{2})+\mathbf{E_{qk}}^{2}\odot(\mathbf{1}_{d}^{\top}\big|\mathbf{Z_{k}}\big|^{2})^{\top}\mathbf{1}_{m}^{\top}-2\mathbf{E_{qk}}\odot\mathrm{Re}(\mathbf{\tilde{Z}}_{\mathbf{q}}^{\dagger}\mathbf{\tilde{Z}}_{\mathbf{k}})$
1exAttention matrix:
$\mathbf{L}=-\log\left(\mathbf{V}^{C}+\big|\mathbf{R}_{\mathbf{qk}}\big|^{2}+\epsilon\right)\odot\mathbf{M}_{\text{causal}},$
@@
1exOutput with value dynamics: $\mathbf{\bar{Z}_{v}}=\bm{\tilde{\Phi}}_{\mathbf{v}}^{+}\odot(\mathbf{\tilde{Z}}_{\mathbf{v}}\mathbf{\hat{A}}^{\dagger})$
@@
Prediction: $\mathbf{Z}_{\mathbf{v}}^{+}=(e^{\alpha\Delta t}e^{\bm{\lambda}_{\Omega}\Delta t})\odot\mathbf{\bar{Z}_{v}}$
Output Projection: $\big(\operatorname{Re}(\mathbf{Z_{p}}),\operatorname{Im}(\mathbf{Z_{p}})\big)\leftarrow\mathcal{L}_{P}\big(\operatorname{Re}(\mathbf{Z}_{\mathbf{v}}^{+}),\operatorname{Im}(\mathbf{Z}_{\mathbf{v}}^{+})\big)$
@@
Algorithm 1 Adaptive Filter Attention (Isotropic; Single Head)
@@ 4.2 Basic Experiments
While standard attention is also able to filter this noisy trajectory, it is not able to do so effectively in a single layer, and the resulting attention matrices lack the consistent and interpretable attention matrices of AFA. We compare the attention matrices from a trained model comprised of two layers of standard attention, with an additive positional embedding, to a trained model comprised of a single layer of AFA in Fig. 3. The AFA attention matrix displays a clear periodic structure, corresponding to pair-wise precisions propagated through the underlying oscillatory dynamics.
@@ 5 Radial-Tangential Model
As we saw in Section 3, we were able to significantly simplify the adaptive filter by assuming that the dynamics decouple into separate magnitude and angular components (i.e. an isotropic plus skew-symmetric state matrix), and that the process and measurement noise are isotropic. Here, we will slightly relax these assumptions by assuming that the process and measurement noise also decompose into separate radial and tangential components, and derive a modification of the above algorithm.
@@ 5.1 Itô SDE with Radial–Tangential Dynamics and Noise
$d\mathbf{x}(t)=\mathbf{A}\mathbf{x}(t)dt+\mathbf{G}(t)d\mathbf{w}(t),\quad\mathbf{z}(t_{k})=\mathbf{C}\mathbf{x}(t_{k})+\mathbf{v}(t_{k}),$
Let $\mathbf{A}(t)=\mathbf{S}\bm{\Lambda}\mathbf{S}^{-1}$, with $\bm{\Lambda}=\alpha\mathbf{I}+\bm{\Lambda}_{\Omega}$, where $\bm{\Lambda}_{\Omega}\in i\mathbb{R}^{d\times d}$ is diagonal, and $\mathbf{G}(t)=\mathbf{S}\bm{\Lambda}_{B}(t)\mathbf{S}^{\dagger}$, $\mathbf{C}(t)=\lambda_{C}\mathbf{I}$. We can also write the eigenvalues as $\bm{\lambda}=\alpha\mathbf{1}+\bm{\omega}$, where $\alpha\in\mathbb{R}$, $\bm{\omega}=\text{diag}(\bm{\Lambda}_{\Omega})\in i\mathbb{R}^{d}$.
@@
$m_{\mathbf{s}}(t)=\big|\mathbf{x}_{\mathbf{s}}(t)\big|_{2}=\sqrt{\mathbf{x}_{\mathbf{s}}^{\dagger}(t)\mathbf{x}_{\mathbf{s}}(t)}\in\mathbb{R},\quad\mathbf{u}_{\mathbf{s}}(t)=\mathbf{x}_{\mathbf{s}}(t)/\big|\mathbf{x}_{\mathbf{s}}(t)\big|_{2}\in\mathbb{S}^{d-1}.$
$m_{\mathbf{zs}}(t)=\big|\mathbf{z}_{\mathbf{s}}(t)\big|_{2}=\sqrt{\mathbf{z}_{\mathbf{s}}^{\dagger}(t)\mathbf{z}_{\mathbf{s}}(t)}\in\mathbb{R},\quad\mathbf{u}_{\mathbf{zs}}(t)=\mathbf{z}_{\mathbf{s}}(t)/\big|\mathbf{z}_{\mathbf{s}}(t)\big|_{2}\in\mathbb{S}^{d-1}.$
@@
The process noise rotates with the dynamics when viewed from the global frame, which corresponds to a fixed anisotropic noise applied in the token’s frame.
@@
$\exp_{\mathbf{u}}:T_{\mathbf{u}}\mathbb{S}^{d-1}\to\mathbb{S}^{d-1},\quad\exp_{\mathbf{u}}(\mathbf{v})=\cos\big(\big|\mathbf{v}\big|\big.)\,\mathbf{u}+\sin\big(\big|\mathbf{v}\big|\big.)\,\frac{\mathbf{v}}{\big|\mathbf{v}\big|},$
@@ 5.2 Propagation of Uncertainty through the RT-SDE
$\mathbf{x}(t_{i})=e^{\mathbf{A}(t_{i}-t_{j})}\mathbf{x}(t_{j})+\int_{t_{j}}^{t_{i}}e^{\mathbf{A}(t_{i}-s)}\mathbf{G}(s)d\mathbf{w}(s)$
@@
$\mathbf{x}(t_{i})=e^{-\mathbf{A}\Delta t_{ij}}\left(\mathbf{x}(t_{j})+\int_{0}^{-\Delta t_{ij}}e^{-\mathbf{A}\tau}\mathbf{G}(t_{j}+\tau)d\mathbf{w}(\tau)\right).$
@@
$\bm{\Lambda}_{\mathbf{V}}^{C}(t_{i},t_{j})=\big|\lambda_{C}\big|^{2}\bm{\Lambda}_{\mathbf{V}F}(t_{i},t_{j})+e^{-\bm{\Lambda}\Delta t_{ij}}\bm{\Lambda}_{R}(t_{j})e^{-\bm{\Lambda}^{\dagger}\Delta t_{ij}}$
$\big|\lambda_{C}\big|^{2}\bm{\Lambda}_{\mathbf{V}F}(t_{i},t_{j})+e^{-\bm{\Lambda}\Delta t_{ij}}\bm{\Lambda}_{R}(t_{j})e^{-\bm{\Lambda}^{\dagger}\Delta t_{ij}}$
$\big|\lambda_{C}\big|^{2}\bm{\Lambda}_{\mathbf{V}F}(t_{i},t_{j})+e^{-2\alpha\Delta t_{ij}}\bm{\Lambda}_{R}(t_{i})$
@@
$\bm{\Lambda}_{\mathbf{V}}^{C}(t_{i},t_{j})$ $=\Big(\big|\lambda_{C}\big|^{2}\kappa_{F}(|\Delta t_{ij}|)\sigma_{t}^{2}+e^{-2\alpha\Delta t_{ij}}\eta_{t}^{2}\Big)\mathbf{I}$
$\quad+\Big(\big|\lambda_{C}\big|^{2}\kappa_{F}(|\Delta t_{ij}|)(\sigma_{r}^{2}-\sigma_{t}^{2})+e^{-2\alpha\Delta t_{ij}}(\eta_{r}^{2}-\eta_{t}^{2})\Big)\mathbf{u_{s}}(t_{i})\mathbf{u_{s}}(t_{i})^{\dagger}$
@@
$\sigma_{Vr}^{2}(|\Delta t_{ij}|)$ $=\big|\lambda_{C}\big|^{2}\,\kappa_{F}(|\Delta t_{ij}|)\,\sigma_{r}^{2}+e^{-2\alpha\Delta t_{ij}}\,\eta_{r}^{2},$
$\sigma_{Vt}^{2}(|\Delta t_{ij}|)$ $=\big|\lambda_{C}\big|^{2}\,\kappa_{F}(|\Delta t_{ij}|)\,\sigma_{t}^{2}+e^{-2\alpha\Delta t_{ij}}\,\eta_{t}^{2}.$
@@
$\bm{\Lambda}_{\mathbf{V}}^{C}(t_{i},t_{j})=\Big(\big|\lambda_{C}\big|^{2}\kappa_{F}(|\Delta t_{ij}|)\sigma^{2}+e^{-2\alpha\Delta t_{ij}}\eta^{2}\Big)\mathbf{I}$
@@ 5.3 Radial–Tangential Kalman IRLS
$m_{\mathbf{zs}i}^{(k)}=\big|\mathbf{z}_{\mathbf{s}i}^{(k)}\big|_{2},\quad\mathbf{u}_{\mathbf{zs}i}^{(k)}=\mathbf{z}_{\mathbf{s}i}^{(k)}/\big|\mathbf{z}_{\mathbf{s}i}^{(k)}\big|_{2}.$
@@
$\begin{gathered}\bm{\Lambda}_{\mathbf{V}}^{C}(t_{i},t_{j})=\sigma_{Vt}^{2}(|\Delta t_{ij}|)\big(\mathbf{I}-\mathbf{u_{s}}(t_{i})\mathbf{u_{s}}(t_{i})^{\dagger}\big)+\sigma_{Vr}^{2}(|\Delta t_{ij}|)\mathbf{u_{s}}(t_{i})\mathbf{u_{s}}(t_{i})^{\dagger},\\[5.69054pt] \sigma_{Vr}^{2}(|\Delta t_{ij}|)=\big|\lambda_{C}\big|^{2}\kappa_{F}(|\Delta t_{ij}|)\sigma_{r}^{2}+e^{-2\alpha\Delta t_{ij}}\eta_{r}^{2},\\[2.84526pt] \sigma_{Vt}^{2}(|\Delta t_{ij}|)=\big|\lambda_{C}\big|^{2}\kappa_{F}(|\Delta t_{ij}|)\sigma_{t}^{2}+e^{-2\alpha\Delta t_{ij}}\eta_{t}^{2}.\end{gathered}$
@@
$\mathbf{P}_{T\mathbf{zs}i}^{(k)}\Big(\mathbf{z}_{\mathbf{s}i}^{(k)}/\lambda_{C}\Big):=\mathbf{I}-\frac{\mathbf{z}_{\mathbf{s}i}^{(k)}\Big(\mathbf{z}_{\mathbf{s}i}^{(k)}\Big)^{\dagger}/\big|\lambda_{C}\big|^{2}}{\Big(\mathbf{z}_{\mathbf{s}i}^{(k)}\Big)^{\dagger}\Big(\mathbf{z}_{\mathbf{s}i}^{(k)}\Big)/\big|\lambda_{C}\big|^{2}}=\mathbf{I}-\mathbf{u}_{\mathbf{zs}i}^{(k)}(\mathbf{u}_{\mathbf{zs}i}^{(k)})^{\dagger}.$
@@
$\theta_{i}^{(k)}:=\big|\mathbf{\bar{v}}_{\mathbf{zs}i}^{(k)}\big|,\qquad\mathbf{a}_{i}^{(k)}:=\frac{\mathbf{\bar{v}}_{\mathbf{zs}i}^{(k)}}{\big|\mathbf{\bar{v}}_{\mathbf{zs}i}^{(k)}\big|},$
@@
$\mathbf{y}_{\mathbf{zs}i}^{(k+1)}=\mathbf{u}_{\mathbf{zs}i}^{(k)}+\mathbf{\bar{v}}_{\mathbf{zs}i}^{(k)}/m_{\mathbf{zs}i}^{(k)},\qquad\mathbf{u}_{\mathbf{zs}i}^{(k+1)}=\frac{\mathbf{y}_{\mathbf{zs}i}^{(k+1)}}{\big|\mathbf{y}_{\mathbf{zs}i}^{(k+1)}\big|},$
@@
$\mathbf{y}_{\mathbf{zs}i}^{(k+1)}=\mathbf{z}_{\mathbf{s}i}^{(k)}+\mathbf{\bar{v}}_{\mathbf{zs}i}^{(k)},\qquad\mathbf{u}_{\mathbf{zs}i}^{(k+1)}=\frac{\mathbf{y}_{\mathbf{zs}i}^{(k+1)}}{\big|\mathbf{y}_{\mathbf{zs}i}^{(k+1)}\big|},$